slipknot3066:

Find the equation of a line, in slope-intercept form of a line that passes through the point (9, 2) and is perpendicular to the line 2x - y = 8.

slipknot3066:

@Aeon

slipknot3066:

@dude

dude:

First you want to turn the equation they gave you into slope-intercept form (Isolate y) Do you know how to do this?

Narad:

The equation of the perpendicular line is \[2x-y=8\] \[y=2x-8\] The slope is \[m _{1}=2\] When 2 lines are perpendicular, the product of their slopes is =-1 \[m _{1}*m _{2}=-1\] Therefore, the slope of the line is \[m _{2}=(-1)/(2)=-1/2\] The equation of the line passing through (9,2) and slope=-1/2 is \[(y-2)=-1/2(x-9)\] \[y=-1/2x+9/2+2=-1/2x+13/2\]